Circular No. 6092

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) Phone 617-495-7244/7440/7444 TWX 710-320-6842 ASTROGRAM CAM

MARSDEN@CFA.HARVARD.EDU or GREEN@CFA.HARVARD.EDU (science)

COMET MACHHOLZ (1994r)

The following precise positions have been reported:

1994 UT  α_{2000}  δ_{2000}  m_{1} Observer
Oct. 9.33938 8^{h}40^{m}07.^{s}41 +55^{o}21′24′′.1 12.5 Viscome
9.34947 8 40 06.18 +55 21 21.4  _{”}
9.35072 8 40 06.07 +55 21 21.5  _{”}
9.36258 8 40 04.80 +55 21 18.4  _{”}
9.44497 8 39 55.64 +55 20 58.8 12.0 Helin
9.48750 8 39 50.99 +55 20 47.1  _{”}

G. Viscome (Lake Placid, NY). 0.37-m f/6 reflector + telecompressor +

CCD. E. Helin, K. Lawrence, B. Arthur and J. Smallwood (Palomar). 0.46-m

Schmidt. Measurer Lawrence. Object slightly diffuse.

Total visual magnitude estimates: Oct. 9.39 UT, 11.3 (A. Hale, near Chaparral, NM, 0.41-m reflector); 9.49, 11–12 (D. E. Machholz, Colfax, CA, 0.25-m reflector).

PERIODIC COMET REINMUTH 2 (1993g) Total magnitude estimates, visual unless otherwise stated: Aug. 7.35

UT, 13.2 (Hale); 14.42, 13.4 (Hale); Sept. 6.98, 14.7 (P. Pravec, Ondřejov, Czech Republic, 0.65-m reflector + CCD); 9.33, 13.5 (Hale); Oct. 5.98, [13.5 (S. Garro, Hautes-Alpes, France, 0.20-m reflector).

SATURN V. J. Tejfel and G. A. Kharitonova, Laboratory of Lunar and Planetary

Physics, Kazakhstan Academy of Sciences, Alma-Ata, report: “During photographic observations with the 1-m telescope at the high-altitude observing station at Assy in good atmospheric conditions during Sept. 24–30 we independently detected a light spot (cf. IAUC 6079, 6080) in Saturn’s equatorial region extending ∼ 30^{o}–35^{o} in longitude. The spot is most clearly visible on photographs in blue and violet and has less contrast in green, orange and red. Its longitude (uncertainty ±2^{o} in both systems I and III changed noticeably: Sept. 25.712 UT, L_{I} = 224^{o}, L_{III} = 56^{o}; 28.733, 253^{o}, 344^{o}; 29.597, 265^{o}, 327^{o}. From these observations the spot’s own rotation period was estimated tentatively as 10^{h}22^{m}.”

1994 October 10 Brian G. Marsden